# Nonreciprocal superradiant quantum phase transition induced by magnon Kerr effect

### v1

@@ Abstract
Recently, proposals for realizing a nonreciprocal superradiant quantum phase transition (SQPT) have been put forward, based on either nonreciprocal interactions between two spin ensembles or the Sagnac-Fizeau shift in a spinning cavity. However, experimental implementation of such a nonreciprocal SQPT remains challenging. This motivates the search for new mechanisms capable of producing a nonreciprocal SQPT. Here, we propose an alternative approach to realize a nonreciprocal SQPT, induced by the magnon Kerr effect (MKE), in a cavity magnonic system, where magnons in a yttrium iron garnet (YIG) sphere are coupled to cavity photons. The Kerr coefficient $K$ is positive (negative), i.e., $K>0$ ($K<0$), when the bias magnetic field is aligned along the crystallographic axis [100] ([110]) of the YIG sphere. We show that the steady-state phase diagrams differ significantly for both cases with $K>0$ and $K<0$, which is the origin of the nonreciprocal SQPT. By further studying the steady-state magnon occupation and its fluctuations versus the parametric drive strength, we demonstrate that the SQPT becomes nonreciprocal, characterized by distinct critical thresholds for $K>0$ and $K<0$. Moreover, we introduce a bidirectional contrast ratio to quantify this nonreciprocal behavior. Our work provides a new mechanism for realizing the nonreciprocal SQPT, with potential applications in designing nonreciprocal quantum devices.
@@ I Introduction
Recently, Fruchart et al. established a general theory of nonreciprocal phase transitions, showing that time-dependent phases arise from asymmetric interactions among multiple species or fields Fruchart21 . The nonreciprocal phenomenon refers to that the system responds differently when an external field is applied in opposite directions. Due to potential applications in quantum technologies, various nonreciprocal quantum effects have been studied, such as thermal noise routers Barzanjeh18 , nonreciprocal photon blockade Huang18 , one-way quantum amplifiers Metelmann15 ; Malz18 ; Shen18 , and nonreciprocal magnon blockade Wang22 . Following the framework of Ref. Fruchart21 , the nonreciprocal SQPT was proposed in a nonreciprocal Dicke model, with quantum-light-mediated nonreciprocal interactions between two spin ensembles Chiacchio23 . Separately, two groups have independently proposed realizing a nonreciprocal SQPT by exploiting the Sagnac-Fizeau shift in a spinning cavity Zhu24 ; Xu24 . However, the key elements of these proposals-specifically, the required nonreciprocal interactions between two spin ensembles Chiacchio23 and the Sagnac-Fizeau shift in a spinning resonator Zhu24 ; Xu24 -present major challenges for existing experimental quantum platforms that have demonstrated the SQPT Baumann10 ; Baumann11 ; Brennecke13 ; Baden14 ; Zhang21Chen ; Wu23 ; Feng15 ; Zheng23 ; Ning24 ; Cai21 ; Zhao25 ; Chen21 ; Wu24 . It is therefore imperative to explore new theoretical mechanisms for realizing a nonreciprocal SQPT.
@@
Owing to their excellent tunability and design flexibility, cavity magnonic systems have emerged as well-suited platforms for exploring novel physical phenomena Soykal10 ; Huebl13 ; Tabuchi14 ; Zhang14 ; Goryachev14 ; Zhang15 ; Zhang15Zou ; Bai15 ; Lachance-Quirion19 ; Rameshti22 ; Yuan22 . In Ref. Wang16 , frequency shifts of both the magnon and cavity modes induced by the magnon Kerr effect (MKE) were experimentally observed by directly driving a yttrium iron garnet (YIG) sphere within a microwave cavity. The MKE, which originates from magnetocrystalline anisotropy in YIG, describes magnon-magnon interactions. It can lead to multistability of cavity magnon polaritons Wang18 ; Zhang19 ; Nair20 ; Bi21 ; Shen21 ; Shen22 , magnon-photon entanglement Zhang19Scully ; Yang21Jin , long-distance spin-spin coupling Xiong22Tian , and unidirectional microwave transmission Kong19 ; Ullah24 ; Miao24 , among other effects. A notable feature of the MKE is that the Kerr coefficient depends on the direction of the bias magnetic field applied to the YIG sphere: the Kerr coefficient is positive when the field is aligned with the crystallographic axis [100], and negative when aligned with the axis [110] Wang18 ; Zhang19 . It is this directional dependence that makes cavity magnonic systems a promising platform for studying nonreciprocal quantum effects, such as nonreciprocal entanglement Chen23 ; Chen24 ; Ahmed24 ; Kong24 , nonreciprocal steering Kong25 , nonreciprocal magnon blockade Zhang24 , and nonreciprocal mechanical squeezing Tong25 .
@@
In this work, we propose a scheme for realizing a nonreciprocal SQPT induced by MKE in a cavity magnonic system. The system consists of a magnon mode in a YIG sphere coupled to a cavity mode via the magnetic-dipole interaction, where the cavity is subject to a parametric drive, and the magnon mode exhibits the MKE. By analyzing the stability of the solutions for the magnon occupation, we obtain the steady-state phase diagram, which comprises three distinct phases: normal phase, superradiant phase and bistable phase. The phase diagram differs significantly when the bias magnetic field is aligned along the [100] or [110] crystallographic axis of the YIG sphere, corresponding to a positive ($K>0$) or negative ($K<0$) Kerr coefficient, respectively Wang18 ; Zhang19 . We further examine the steady-state magnon number and its fluctuations as functions of the parametric drive strength, revealing that the SQPT becomes nonreciprocal, with different critical thresholds for $K>0$ and $K<0$. To quantify this behavior, we introduce a bidirectional contrast ratio that characterizes the degree of nonreciprocity in the SQPT. Our work provides an alternative approach for realizing a nonreciprocal SQPT Chiacchio23 ; Zhu24 ; Xu24 and offers useful guidance for the experimental realization of this phenomenon in the future.
@@
The remainder of this paper is structured as follows. In Sec. II, we introduce the proposed cavity magnonic system and derive the steady-state solutions for the magnon occupation. Section III presents the steady-state phase diagram based on a stability analysis of these solutions, followed by a detailed investigation of the MKE-induced nonreciprocal SQPT. Finally, a brief summary is given in Sec. IV.
@@ II Model
As depicted in Fig. 1, the considered cavity magnonic system consists of a microwave cavity and a millimeter-sized YIG sphere, where the cavity is pumped by a parametric drive, and the YIG sphere is magnetized to saturation by a bias static magnetic field. The YIG sphere supports a Kittle mode (with frequency $\omega_{m}$), which is coupled to the cavity mode (with frequency $\omega_{a}$) via the collective magnetic-dipole interaction. Taking into account both the system’s dissipation and the interaction among magnons (i.e., MKE), the parametrically driven cavity magnonic system can be described by the following effective non-Hermitian Hamiltonian (see Appendix A):
@@
where $a^{{\dagger}}$ and $a$ ( $m^{{\dagger}}$ and $m$) are the creation and annihilation operators of the cavity mode (magnon mode), $\Delta_{a(m)}=\omega_{a(m)}-\omega_{d}/2$ $(>0)$ is the frequency detuning of the cavity mode (magnon mode) from the drive field, $\kappa_{a}$ ($\gamma_{m}$) is the decay rate of the cavity mode (magnon mode), $K$ is the Kerr coefficient, $g_{m}$ is the coupling strength between cavity mode and magnon mode, and $\omega_{d}$ ($\Omega$) is the frequency (strength) of parametric drive. Here the MKE results from the magnetocrystalline anisotropy of the YIG sphere, and the sign of Kerr coefficient $K$ can be controlled via adjusting the direction of the bias magnetic field Wang18 ; Zhang19 . When the bias magnetic field is aligned along the crystallographic axis [100], the Kerr coefficient $K$ is positive ($K>0$). In contrast, $K$ becomes negative ($K<0$) for aligning the bias magnetic field along the crystallographic axis [110]. The dependence of $K$ on the direction of the bias magnetic field is the origin of the nonreciprocal SQPT (cf. Sec. III). The effective Hamiltonian $H_{\rm eff}$ maintains invariance under the transformation $a\rightarrow-a$ and $b\rightarrow-b$, demonstrating parity symmetry Emary03PRL ; Emary03PRE . Defining the parity operator $\prod=\exp[i\pi(a^{\dagger}a+b^{\dagger}b)]$, this symmetry can be expressed as the commutation relation $H_{\rm eff},\prod]=0$. In our system, the spontaneous breaking of parity symmetry occurs when the cavity magnonic system undergoes a phase transition from the normal phase to superradiant phase.
@@
$|M|^{2}_{+}>0$ $\Omega>\Omega_{1}$ for $\Delta_{m}/\Delta_{a}<\xi$ or $\Omega>\Omega_{2}$ for $\Delta_{m}/\Delta_{a}>\xi$ trivial solution
$|M|^{2}_{-}>0$ trivial solution $\Omega>\Omega_{2}$ for $\Delta_{m}/\Delta_{a}<\xi$ or $\Omega>\Omega_{1}$ for $\Delta_{m}/\Delta_{a}>\xi$
@@
$|M|^{2}_{0}=0,~~~|M|^{2}_{\pm}=\Big(-\Delta^{\prime}_{m}\pm\sqrt{\eta^{2}G^{2}-{\gamma^{\prime}_{m}}^{2}}\Big)/K,$ (5)
@@
In this study, we choose the magnon number $|M|^{2}$ as the order parameter for the SQPT. The nonnegativity constraint on the magnon number $|M|^{2}$ restricts the valid solutions to $|M|^{2}_{\pm}\geq 0$. Table 1 summarizes the conditions for $|M|^{2}_{+}>0$ when $K>0$ and $|M|^{2}_{-}>0$ when $K<0$, where we exclude the trivial solutions $|M|^{2}_{-}$ when $K>0$ and $|M|^{2}_{+}$ when $K<0$ due to their instability across all parameters (see Fig. 2 and related discussions). For $K>0$, the constraint $|M|^{2}_{+}>0$ requires $\Omega>\Omega_{1}$ when $\Delta_{m}/\Delta_{a}<\xi$ or $\Omega>\Omega_{2}$ for $\Delta_{m}/\Delta_{a}>\xi$, with the critical ratio $\xi$ defined as
@@
which determined the boundaries between different phases in the steady-state phase diagram of the cavity magnonic system (see Fig. 2). In particular, $\Omega_{1}=\Omega_{2}$ at $\Delta_{m}/\Delta_{a}=\xi$. For $K<0$, the constraint $|M|^{2}_{-}>0$ is satisfied when $\Omega>\Omega_{2}$ for $\Delta_{m}/\Delta_{a}<\xi$ or $\Omega>\Omega_{1}$ for $\Delta_{m}/\Delta_{a}>\xi$.
@@ III Nonreciprocal superradiant quantum phase transition
bistable phase $|M|^{2}=0$ and $|M|^{2}=|M|^{2}_{+}$ $|M|^{2}=0$ and $|M|^{2}=|M|^{2}_{-}$
Based on the stability analysis of the solutions in Eq. (5), we plot the steady-state phase diagram of the system as a function of the normalized drive strength $\Omega/\kappa_{a}$ and the detuning ratio $\Delta_{m}/\Delta_{a}$ in Fig. 2. Note that the white background in the upper-right corner of Fig. 2(b) indicates the unstable regime, where all three solutions for $|M|^{2}$ are unstable. Figures 2(a) and 2(b) correspond to $K>0$ and $K<0$, respectively, which can be realized by aligning the bias magnetic field parallel to the [100] and [110] crystallographic axes of the YIG sphere Wang18 ; Zhang19 . The difference between these two panels demonstrates the nonreciprocity of the SQPT. The phase diagrams exhibits three distinct phases: normal phase (light blue area), superradiant phase (light green area) and bistable phase (light orange area), where the boundaries among them are determined by the vertical blue solid line $\Omega=\Omega_{1}$ given in Eq. (8)] and the blue dash-dotted curve $\Omega=\Omega_{2}$ given in Eq. (9)]. Table 2 summarizes the stable solutions in each phase. In the normal phase, the solution $|M|^{2}=0$ is stable for both $K>0$ and $K<0$, corresponding to the vacuum state. In the superradiant phase, the solution $|M|^{2}=|M|^{2}_{+}$ for $K>0$ ($|M|^{2}=|M|^{2}_{-}$ for $K<0$) becomes stable, and the system exhibits macroscopic magnon excitations. In the bistable phase, both $|M|^{2}=0$ and $|M|^{2}=|M|^{2}_{+}$ for $K>0$ ($|M|^{2}=0$ and $|M|^{2}=|M|^{2}_{-}$ for $K<0$) are stable, and the system’s initial state determines whether macroscopic magnon excitations occur Zhang21-Chen . For the parameters used in Fig. 2, the critical ratio $\xi$ from Eq. (7) is $\xi=0.976$. When $\Delta_{m}/\Delta_{a}<\xi$, the boundary between normal phase and bistable phase (normal phase and superradiant phase) is given by $\Omega=\Omega_{1}$ ($\Omega=\Omega_{2}$) for $K>0$ ($K<0$). Conversely, when $\Delta_{m}/\Delta_{a}>\xi$, the boundary between normal phase and superradiant phase (normal phase and bistable phase) is given by $\Omega=\Omega_{2}$ ($\Omega=\Omega_{1}$) for $K>0$ ($K<0$). Regardless of the value of $\Delta_{m}/\Delta_{a}$ and the sign of $K$, the curve $\Omega=\Omega_{2}$ universally defines the boundary between bistable phase and superradiant phase.
@@
In Figs. 3(a), we displays the scaled steady-state magnon number $|M|^{2}/|(\gamma_{m}/K)|$ as a function of the normalized drive strength $\Omega/\kappa_{a}$ for a fixed detuning ratio $\Delta_{m}/\Delta_{a}=1.3$ ($\xi=0.976$). Yellow shaded regions indicate the nonreciprocal domains of SQPT, where $|M|^{2}(K>0)\neq|M|^{2}(K<0)$. In the case of $K<0$, the magnon number exhibits a discontinuous jump from zero to a finite value at the critical point $\Omega/\kappa_{a}=2.025$ as the normalized drive strength $\Omega/\kappa_{a}$ increases across this threshold (blue dashed curve), signifying a second-order SQPT from the normal phase to the superradiant phase. However, for $K>0$, the transition becomes first-order (i.e., a continuous SQPT), and the critical point shifts to $\Omega/\kappa_{a}=2.108$ (red solid curve). Figure 3(b) shows $|M|^{2}/|(\gamma_{m}/K)|$ versus $\Omega/\kappa_{a}$ by taking $\Delta_{m}/\Delta_{a}=0.8$ ($\xi=0.976$), where nonreciprocal SQPT is also evident. Unlike the case with $\Delta_{m}/\Delta_{a}>\xi$ [cf. Figs. 3(a) and 3(b)], the system now undergoes a second-order SQPT from the normal phase to the superradiant phase at $\Omega/\kappa_{a}=2.084$ for $K<0$ (blue dashed curve), while the transition remains first-order at $\Omega/\kappa_{a}=2.025$ if $K>0$ (red solid curve).
In order to quantitatively characterize the nonreciprocity in the SQPT, we introduce the bidirectional contrast ratio:
which satisfies $0\leq\mathcal{I}\leq 1$. This parameter quantifies the degree of nonreciprocity in the SQPT: a nonzero $\mathcal{I}$ signals the emergence of nonreciprocal behavior. The minimum value $\mathcal{I}_{\rm min}=0$ corresponds to a reciprocal SQPT, while the maximum value $\mathcal{I}_{\rm max}=1$ represents an ideal nonreciprocal SQPT. In Fig. 4, we show the bidirectional contrast ratio $\mathcal{I}$ as a function of the normalized drive strength $\Omega/\kappa_{a}$ and the detuning ratio $\Delta_{m}/\Delta_{a}$. The white region in the upper-right corner corresponds to the unstable regime identified in Fig. 2(b). The vertical blue solid line (red dashed-dotted curve) marks the critical drive strength $\Omega=\Omega_{1}=2.025$ ($\Omega=\Omega_{2}$). The bidirectional contrast ratio $\mathcal{I}$ attains its minimum value $\mathcal{I}_{\rm min}=0$ in the black region ($\Omega<\Omega_{1}$), where the system remains in the normal phase for both $K>0$ and $K<0$, indicating reciprocal SQPT. Notably, two red regions appear for $\Omega_{1}<\Omega<\Omega_{2}$, where $\mathcal{I}$ reaches its maximum value $\mathcal{I}_{\rm max}=1$, corresponding to an ideal nonreciprocal SQPT. In the lower (upper) red region, the system is in the superradiant phase (normal phase) for $K>0$ but remains in the normal phase (superradiant phase) for $K<0$. In the remaining region ($\Omega>\Omega_{2}$), the bidirectional contrast ratio lies between 0 and 1 (i.e., $0<\mathcal{I}<1$). Here, the system is in the superradiant phase for both $K>0$ and $K<0$, but the order parameters differ: $|M|^{2}(K>0)=|M|^{2}_{+}\neq|M|^{2}(K<0)=|M|^{2}_{-}$.
Figure 4: Bidirectional contrast ratio $\mathcal{I}$ as a function of the normalized drive strength $\Omega/\kappa_{a}$ and the detuning ratio $\Delta_{m}/\Delta_{a}$. The white background in the upper-right corner corresponds to the unstable region in Fig. 2(b). The vertical blue solid line marks the critical drive $\Omega=\Omega_{1}$, and the red dashed-dotted curve indicates $\Omega=\Omega_{2}$. Other parameters are the same as those in Fig. 2.

### v2

@@ Abstract
Recently, proposals for realizing a nonreciprocal superradiant quantum phase transition (SQPT) have been put forward, based on either nonreciprocal interactions between two spin ensembles or the Sagnac-Fizeau shift in a spinning cavity. However, experimental implementation of such a nonreciprocal SQPT remains challenging. This motivates the search for new mechanisms capable of producing a nonreciprocal SQPT. Here, we propose an alternative approach to realize a nonreciprocal SQPT, induced by the magnon Kerr effect (MKE), in a cavity magnonic system, where magnons in a yttrium iron garnet (YIG) sphere are coupled to cavity photons. The MKE coefficient is positive ($K>0$) when the bias magnetic field is aligned along the crystallographic axis [100], but negative ($K<0$) when aligned along the axis [110]. We show that the steady-state phase diagram for $K>0$ differs markedly from that for $K<0$. This contrast is the origin of the nonreciprocal SQPT. By further studying the steady-state magnon occupation and its fluctuations versus the parametric drive strength, we demonstrate that the SQPT becomes nonreciprocal, characterized by distinct critical thresholds for $K>0$ and $K<0$. Moreover, we introduce a bidirectional contrast ratio to quantify this nonreciprocal behavior. Our work provides a new mechanism for realizing the nonreciprocal SQPT, with potential applications in designing nonreciprocal quantum devices.
@@ I Introduction
Recently, Fruchart et al. established a general theory of nonreciprocal phase transitions, showing that time-dependent phases arise from asymmetric interactions among multiple species or fields Fruchart21 . The nonreciprocal phenomenon refers to that the system responds differently when an external field is applied in opposite directions Barzanjeh25 . Due to potential applications in quantum technologies, various nonreciprocal quantum effects have been studied, such as thermal noise routers Barzanjeh18 , one-way quantum amplifiers Metelmann15 ; Malz18 ; Shen18 , and nonreciprocal photon (or magnon) blockade Huang18 ; Wang22 . Following the framework of Ref. Fruchart21 , the nonreciprocal SQPT was proposed in a nonreciprocal Dicke model, with quantum-light-mediated nonreciprocal interactions between two spin ensembles Chiacchio23 . In a parallel approach, two groups have independently proposed realizing a nonreciprocal SQPT by exploiting the Sagnac-Fizeau shift in a spinning cavity Zhu24 ; Xu24 . However, the key elements of these proposals—specifically, the required nonreciprocal interactions between two spin ensembles Chiacchio23 and the Sagnac-Fizeau shift in a spinning cavity Zhu24 ; Xu24 —present major challenges for existing experimental quantum platforms that have demonstrated the SQPT Baumann10 ; Baumann11 ; Brennecke13 ; Baden14 ; Zhang21Chen ; Wu23 ; Feng15 ; Zheng23 ; Ning24 ; Cai21 ; Zhao25 ; Chen21 ; Wu24 . It is therefore imperative to explore new theoretical mechanisms for realizing a nonreciprocal SQPT.
@@
Owing to their excellent tunability and design flexibility, cavity magnonic systems have emerged as well-suited platforms for exploring novel physical phenomena Soykal10 ; Huebl13 ; Tabuchi14 ; Zhang14 ; Goryachev14 ; Zhang15 ; Zhang15Zou ; Bai15 ; Lachance-Quirion19 ; Rameshti22 ; Yuan22 . In Ref. Wang16 , frequency shifts of both the magnon and cavity modes induced by the magnon Kerr effect (MKE) were experimentally observed by directly driving a yttrium iron garnet (YIG) sphere within a microwave cavity. The MKE, originating from magnetocrystalline anisotropy in YIG, describes magnon-magnon interactions. It can lead to multistability of cavity magnon polaritons Wang18 ; Zhang19 ; Nair20 ; Bi21 ; Shen21 ; Shen22 , magnon-photon entanglement Zhang19Scully ; Yang21Jin , long-distance spin-spin coupling Xiong22Tian , and unidirectional microwave transmission Kong19 ; Ullah24 ; Miao24 , among other effects. A notable feature of the MKE is that its coefficient depends on the direction of the bias magnetic field applied to the YIG sphere: the MKE coefficient is positive when the field is aligned with the crystallographic axis [100], but negative when aligned with the axis [110] Wang18 ; Zhang19 . It is this directional dependence that makes cavity magnonic systems a promising platform for studying nonreciprocal quantum effects, such as nonreciprocal quantum entanglement Chen23 ; Chen24 ; Ahmed24 ; Kong24 ; Liu25 ; Kong25 , nonreciprocal magnon (or photon) blockade Zhang24 ; Fan24 , and nonreciprocal quantum synchronization Lai25 .
@@
In this work, we propose a scheme to realize a nonreciprocal SQPT induced by MKE in a cavity magnonic system. The system consists of a magnon mode in a YIG sphere coupled to a cavity mode via the magnetic-dipole interaction, where the cavity is subject to a parametric drive, and the magnon mode exhibits the MKE. By analyzing the stability of the magnon occupation solutions, we obtain the steady-state phase diagram, which comprises three distinct phases: normal phase, superradiant phase and bistable phase. The phase diagram differs significantly when the bias magnetic field is aligned along the [100] and [110] crystallographic axes of the YIG sphere, corresponding to positive ($K>0$) and negative ($K<0$) MKE coefficients, respectively Wang18 ; Zhang19 . We further examine the steady-state magnon number and its fluctuations as functions of the parametric drive strength, revealing that the SQPT becomes nonreciprocal, with different critical thresholds for $K>0$ and $K<0$. To quantify this behavior, we introduce a bidirectional contrast ratio that characterizes the degree of nonreciprocity in the SQPT. Our work provides an alternative approach for realizing a nonreciprocal SQPT Chiacchio23 ; Zhu24 ; Xu24 .
@@
The remainder of this paper is structured as follows. In Sec. II, we introduce the proposed cavity magnonic system and derive the steady-state solutions for the magnon occupation. Section III presents the steady-state phase diagram based on a stability analysis of these solutions, followed by a detailed investigation of the MKE-induced nonreciprocal SQPT. The feasibility of experimentally implementing the proposed scheme is discussed in Sec. IV, which also provides a brief summary.
@@ II Model
As depicted in Fig. 1, the considered cavity magnonic system consists of a microwave cavity and a millimeter-sized YIG sphere, where the cavity is pumped by a parametric drive, and the YIG sphere is magnetized to saturation by a bias static magnetic field. The YIG sphere supports a Kittel mode (with frequency $\omega_{m}$), which is coupled to the cavity mode (with frequency $\omega_{a}$) via the collective magnetic-dipole interaction. Taking into account both the system’s dissipation and the interaction among magnons (i.e., MKE), the parametrically driven cavity magnonic system can be described by the following effective non-Hermitian Hamiltonian (see Appendix A; hereafter setting $\hbar=1$):
@@
where $a^{{\dagger}}$ and $a$ ( $m^{{\dagger}}$ and $m$) are the creation and annihilation operators of the cavity mode (magnon mode), $\Delta_{a(m)}=\omega_{a(m)}-\omega_{d}/2$ $(>0)$ is the frequency detuning of the cavity mode (magnon mode) from the drive field, $\kappa_{a}$ ($\gamma_{m}$) is the decay rate of the cavity mode (magnon mode), $K$ is the MKE coefficient, $g_{m}$ is the coupling strength between cavity mode and magnon mode, and $\omega_{d}$ ($\Omega$) is the frequency (strength) of parametric drive. Here the MKE results from the magnetocrystalline anisotropy of the YIG sphere, and the sign of MKE coefficient $K$ can be controlled by adjusting the direction of the bias magnetic field Wang18 ; Zhang19 . When the bias magnetic field is aligned along the crystallographic axis [100], the MKE coefficient $K$ is positive ($K>0$). In contrast, $K$ becomes negative ($K<0$) for aligning the bias magnetic field along the crystallographic axis [110]. The dependence of $K$ on the direction of the bias magnetic field is the origin of the nonreciprocal SQPT (cf. Sec. III). The effective Hamiltonian $H_{\rm eff}$ maintains invariance under the transformation $a\rightarrow-a$ and $m\rightarrow-m$, demonstrating parity symmetry Emary03PRL ; Emary03PRE . Defining the parity operator $\Pi=\exp[i\pi(a^{\dagger}a+m^{\dagger}m)]$, this symmetry can be expressed as the commutation relation $H_{\rm eff},\Pi]=0$. In our system, the spontaneous breaking of parity symmetry occurs when the cavity magnonic system undergoes a phase transition from the normal phase to superradiant phase.
@@
$|M|^{2}_{+}>0$ $\Omega>\Omega_{1}$ for $\Delta_{m}/\Delta_{a}<\xi$; $\Omega>\Omega_{2}$ for $\Delta_{m}/\Delta_{a}>\xi$ trivial
$|M|^{2}_{-}>0$ trivial $\Omega>\Omega_{2}$ for $\Delta_{m}/\Delta_{a}<\xi$; $\Omega>\Omega_{1}$ for $\Delta_{m}/\Delta_{a}>\xi$
@@
$|M|^{2}_{0}=0,~~~|M|^{2}_{\pm}=\Big(-\Delta^{\prime}_{m}\pm\sqrt{\eta^{2}\Omega^{2}-{\gamma^{\prime}_{m}}^{2}}\Big)/K,$ (5)
@@
In this study, we choose the magnon number $|M|^{2}$ as the order parameter for the SQPT. The nonnegativity constraint on the magnon number $|M|^{2}$ restricts the valid solutions to $|M|^{2}_{\pm}\geq 0$. Table 1 summarizes the conditions for $|M|^{2}_{+}>0$ when $K>0$ and $|M|^{2}_{-}>0$ when $K<0$, where we exclude the trivial solutions $|M|^{2}=|M|^{2}_{-}$ when $K>0$ and $|M|^{2}=|M|^{2}_{+}$ when $K<0$ due to their instability across all parameters (see Fig. 2 and related discussions). For $K>0$, the constraint $|M|^{2}_{+}>0$ requires $\Omega>\Omega_{1}$ when $\Delta_{m}/\Delta_{a}<\xi$ or $\Omega>\Omega_{2}$ for $\Delta_{m}/\Delta_{a}>\xi$, with the critical ratio $\xi$ defined as
@@
which determine the boundaries between different phases in the steady-state phase diagram of the cavity magnonic system (see Fig. 2). In particular, $\Omega_{1}=\Omega_{2}$ at $\Delta_{m}/\Delta_{a}=\xi$. For $K<0$, the constraint $|M|^{2}_{-}>0$ is satisfied when $\Omega>\Omega_{2}$ for $\Delta_{m}/\Delta_{a}<\xi$ or $\Omega>\Omega_{1}$ for $\Delta_{m}/\Delta_{a}>\xi$.
@@ III Nonreciprocal superradiant quantum phase transition
bistable phase $|M|^{2}=0,\,|M|^{2}_{+}$ $|M|^{2}=0,\,|M|^{2}_{-}$
Based on the stability analysis of the solutions in Eq. (5), we plot the steady-state phase diagram of the system as a function of the normalized drive strength $\Omega/\kappa_{a}$ and the detuning ratio $\Delta_{m}/\Delta_{a}$ in Fig. 2. Note that the white background in the upper-right corner of Fig. 2(b) indicates the unstable regime, where all three solutions for $|M|^{2}$ are unstable. Figures 2(a) and 2(b) correspond to $K>0$ and $K<0$, respectively, which can be realized by aligning the bias magnetic field parallel to the [100] and [110] crystallographic axes of the YIG sphere Wang18 ; Zhang19 . The difference between these two panels demonstrates the nonreciprocity of the SQPT. The phase diagram exhibits three distinct phases: normal phase (light blue area), superradiant phase (light green area) and bistable phase (light orange area), where the boundaries among them are determined by the vertical blue solid line $\Omega=\Omega_{1}$ given in Eq. (8)] and the green dash-dotted curve $\Omega=\Omega_{2}$ given in Eq. (9)]. Table 2 summarizes the stable solutions in each phase. In the normal phase, the solution $|M|^{2}=0$ is stable for both $K>0$ and $K<0$, corresponding to the vacuum state. In the superradiant phase, the solution $|M|^{2}=|M|^{2}_{+}$ for $K>0$ ($|M|^{2}=|M|^{2}_{-}$ for $K<0$) becomes stable, and the system exhibits macroscopic magnon excitations. In the bistable phase, both $|M|^{2}=0$ and $|M|^{2}=|M|^{2}_{+}$ for $K>0$ ($|M|^{2}=0$ and $|M|^{2}=|M|^{2}_{-}$ for $K<0$) are stable, and the system’s initial state determines whether macroscopic magnon excitations occur Zhang21-Chen . For the parameters used in Fig. 2, the critical ratio $\xi$ from Eq. (7) is $\xi=0.976$. When $\Delta_{m}/\Delta_{a}<\xi$, the boundary between normal phase and bistable phase (normal phase and superradiant phase) is given by $\Omega=\Omega_{1}$ ($\Omega=\Omega_{2}$) for $K>0$ ($K<0$). Conversely, when $\Delta_{m}/\Delta_{a}>\xi$, the boundary between normal phase and superradiant phase (normal phase and bistable phase) is given by $\Omega=\Omega_{2}$ ($\Omega=\Omega_{1}$) for $K>0$ ($K<0$). Regardless of the value of $\Delta_{m}/\Delta_{a}$ and the sign of $K$, the curve $\Omega=\Omega_{2}$ universally defines the boundary between bistable phase and superradiant phase.
@@
In Figs. 3(a), we display the scaled steady-state magnon number $|M|^{2}/|(\gamma_{m}/K)|$ as a function of the normalized drive strength $\Omega/\kappa_{a}$ for a fixed detuning ratio $\Delta_{m}/\Delta_{a}=1.3$ ($\xi=0.976$). Yellow shaded regions indicate the nonreciprocal domains of SQPT, where $|M|^{2}(K>0)\neq|M|^{2}(K<0)$. In the case of $K<0$, the magnon number exhibits a discontinuous jump from zero to a finite value at the critical point $\Omega/\kappa_{a}=2.025$ as the normalized drive strength $\Omega/\kappa_{a}$ increases across this threshold (blue dashed curve), signifying a first-order SQPT from the normal phase to the superradiant phase. However, for $K>0$, the transition becomes second-order (i.e., a continuous SQPT), and the critical point shifts to $\Omega/\kappa_{a}=2.108$ (red solid curve). Figure 3(b) shows $|M|^{2}/|(\gamma_{m}/K)|$ versus $\Omega/\kappa_{a}$ by taking $\Delta_{m}/\Delta_{a}=0.8$ ($\xi=0.976$), where nonreciprocal SQPT is also evident. Unlike the case with $\Delta_{m}/\Delta_{a}>\xi$ [cf. Figs. 3(b) and 3(a)], the system now undergoes a second-order SQPT from the normal phase to the superradiant phase at $\Omega/\kappa_{a}=2.084$ for $K<0$ (blue dashed curve), while the transition remains first-order at $\Omega/\kappa_{a}=2.025$ if $K>0$ (red solid curve).